# Learning to Pursue AC Optimal Power Flow Solutions with Feasibility Guarantees

### v1

@@ V Theoretical Analysis
where $\underline{V}_{e}=\underline{V}-\epsilon_{v}-2E_{v}-\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$, $\overline{V}_{e}=\overline{V}+\epsilon_{v}+2E_{v}+\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$, and $\overline{I}_{e}=\overline{I}+\epsilon_{i}+2E_{i}+\|\boldsymbol{\Gamma}_{i}\|\epsilon^{\textsf{NN}}$. Then, the neural network-based algorithm (8) renders a set $\mathcal{S}_{s}$, with $\mathcal{S}\subseteq\mathcal{S}_{s}\subseteq\mathcal{S}_{e}$, forward invariant. $\triangle$
We note that $\mathcal{S}_{s}$ in Proposition V.3, is an inflation of the set of feasible voltages and currents $\mathcal{S}$ specified in the AC OPF. Hence the terminology “practical feasibility”. Indeed, when these errors are small, the constraint violation is practically negligible. We provide the following remarks:
  • i)
    When $\epsilon_{v}$, $\epsilon_{i}$, $E_{v}$, $E_{i}$, and $\epsilon^{\textsf{NN}}$ are available or are estimated numerically, the constraints of the original AC OPF (II-B) can be tightened so that (8) can render the feasible set $\mathcal{S}$ forward invariant.
  • ii)
    If $\boldsymbol{u}(t_{0})\in\mathcal{S}_{s}$, then $\boldsymbol{u}(t)\in\mathcal{S}_{s}$ for all $t\geq t_{0}$. This implies that the solution offered by (8) is practically feasible for both the online implementation in Figure 1(left) and when the offline procedure in Figure 1(center) is terminated before convergence.
  • iii)
    Since $\|\boldsymbol{\Gamma}_{v}\|$ and $\|\boldsymbol{\Gamma}_{i}\|$ are generally small (less than $0.0982$ in our numerical experiments), the constraint violation due to the neural network approximation is practically negligible.
@@ VI Conclusions
We have proposed a solution method for solving the AC OPF where a neural network is used to approximate the solution of a convex QP defining the safe gradient flow. Our approach is shown to lead to both feedback-based online implementations and offline solutions based on power flow computations. Compared to existing methods that rely on neural networks, our algorithm ensures that the DERs’ setpoints are practically feasible and it ensures convergence to a neighborhood of a strict local optimizer of the AC OPF. These guarantees are important for power systems optimization tasks, as operating limits must be satisfied for a safe power delivery.
@@
Recall that $\boldsymbol{\nu}:=|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|+\boldsymbol{n}_{v}$, $\boldsymbol{i}:=|\boldsymbol{i}(\boldsymbol{u};\boldsymbol{s}_{l})|+\boldsymbol{n}_{i}$, $\boldsymbol{\xi}=(\boldsymbol{\nu},\boldsymbol{\iota})$; to streamline notation, we will use $|\boldsymbol{v}|$ and $|\boldsymbol{i}|$ to denote the error-free measurements or computations of voltage magnitudes currents magnitudes. Recall that $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})=F_{\textsf{m}}(\boldsymbol{u},\boldsymbol{0},\boldsymbol{0},\boldsymbol{0})$. We express the NN-SGF controller as $\dot{\boldsymbol{u}}=\mathcal{F}^{\textsf{NN}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$, where $\boldsymbol{\theta}=(\boldsymbol{\theta}_{u,1},\ldots,\boldsymbol{\theta}_{u,G},\underline{V},\overline{V},\overline{I})$ contains the constraint parameters of the AC OPF. Rewrite the NN-SGF controller as:
@@
and is negative definite. Let $e_{1}:=-\lambda_{\max}(\boldsymbol{J}_{F})$, $e_{2}:=-\lambda_{\min}(\boldsymbol{J}_{F})$, and define the matrix
[…]
@@
-B Proof of Theorem V.3
The proof leverages Nagumo’s Theorem. Consider the SGF controller $F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ in (11); let $\hat{v}_{j}:=|\boldsymbol{\Gamma}_{v,j}\boldsymbol{u}+\bar{v}_{j}(\boldsymbol{s}_{l})|$ denote the linearized voltage magnitude at index $j\in\mathcal{M}$, and let the measurement noise $\boldsymbol{n}_{v}$ satisfy $\|\boldsymbol{n}_{v}\|\leq\epsilon_{v}$, as in Assumption 4. Then, for each $j\in\mathcal{M}$:
$-\boldsymbol{\Gamma}_{v,j}^{\top}F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ $\leq-\beta\left(\underline{V}-|\hat{v}_{j}|-\epsilon_{v}-E_{v}\right),$
$\boldsymbol{\Gamma}_{v,j}^{\top}F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ $\leq-\beta\left(|\hat{v}_{j}|-(\overline{V}+\epsilon_{v}+E_{v})\right),$
where $E_{v}$ bounds the voltage linearization error. Now consider the NN-SGF:
@@
$\leq-\beta(\underline{V}-|\hat{v}_{j}|-\epsilon_{v}-E_{v})+\|\boldsymbol{\Gamma}_{v,j}\|\cdot\epsilon^{\textsf{NN}},$
@@
$\leq-\beta(|\hat{v}_{j}|-\overline{V}-\epsilon_{v}-E_{v})+\|\boldsymbol{\Gamma}_{v,j}\|\cdot\epsilon^{\textsf{NN}}.$
To ensure the vector field is inward-pointing at the boundary of the voltage constraint set, we then require:
@@
$\mathcal{S}_{s,\hat{v}}$ :=$:=\left\{\boldsymbol{u}\in\mathcal{C}:\underline{V}_{s}\leq|\hat{\boldsymbol{v}}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V}_{s}\right\},$
$\mathcal{S}_{s,\hat{\imath}}$ :=$:=\left\{\boldsymbol{u}\in\mathcal{C}:|\hat{\boldsymbol{i}}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I}_{s}\right\},$
@@
Since the NN-SGF vector field is strictly inward-pointing, Nagumo’s Theorem implies that $\mathcal{S}_{s}$ is forward invariant under (8). To conclude the proof, we note that $|\hat{v}_{j}|-E_{v}\leq|v_{j}|\leq|\hat{v}_{j}|+E_{v}$ and $|\hat{\imath}_{j}|\leq|\imath_{j}|+E_{i}$, and thus $\mathcal{S}_{s}\subseteq\mathcal{S}_{e}$.

### v2

@@ V Theoretical Analysis
where $\underline{V}_{e}=\underline{V}-\epsilon_{v}-2E_{v}-\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$, $\overline{V}_{e}=\overline{V}+\epsilon_{v}+2E_{v}+\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$, and $\overline{I}_{e}=\overline{I}+\epsilon_{i}+2E_{i}+\beta^{-1}\|\boldsymbol{\Gamma}_{i}\|\epsilon^{\textsf{NN}}$. Then, the neural network-based algorithm (8) renders a set $\mathcal{S}_{s}$, with $\mathcal{S}\subseteq\mathcal{S}_{s}\subseteq\mathcal{S}_{e}$, forward invariant. $\triangle$
We recall that $E_{v}$ and $E_{i}$ are defined in Assumption 3. The set $\mathcal{S}_{s}$ is an auxiliary set defined through linearized voltage and current magnitudes and is instrumental for establishing forward invariance of the NN-SGF dynamics. The formal definition of $\mathcal{S}_{s}$ is provided in the appendix. The set $\mathcal{S}_{e}$, instead, is expressed in terms of the actual voltage and current magnitudes and contains $\mathcal{S}_{s}$ by accounting for the linearization and neural-network approximation errors. This construction leads to the factor $2$ appearing in the definitions of $\underline{V}_{e}$, $\overline{V}_{e}$, and $\overline{I}_{e}$. We note that $\mathcal{S}_{s}$ in Proposition V.3 is an inflation of the set of feasible voltages and currents $\mathcal{S}$ specified in the AC OPF. Hence the terminology “practical feasibility”. Indeed, when these errors are small, the constraint violation is practically negligible. The proofs of all results are provided in the appendix.
We provide the following remarks.
Remark V.4 (Practical forward invariance).
If $\boldsymbol{u}(t_{0})\in\mathcal{S}_{s}$, then $\boldsymbol{u}(t)\in\mathcal{S}_{s}$ for all $t\geq t_{0}$. Consequently, the control input generated by (8) is practically feasible both for the online implementation shown in Fig. 1(left) and when the offline procedure in Fig. 1(center) is terminated prior to convergence. Moreover, the bounds $\underline{V}_{e}$, $\overline{V}_{e}$, and $\overline{I}_{e}$ provide an explicit quantification of the sets of voltage magnitudes and currents that are achievable under (8) as a function of the bounds $\epsilon_{v}$, $\epsilon_{i}$, $E_{v}$, $E_{i}$, and the neural-network approximation error $\epsilon^{\textsf{NN}}$. For instance, in the experiments reported in Section IV, the maximum value of $\|\boldsymbol{\Gamma}_{v}\|$ and $\|\boldsymbol{\Gamma}_{i}\|$ was $0.0982$, $E_{v}\approx 10^{-2}$ p.u. and $\epsilon^{\textsf{NN}}\approx 1.7\times 10^{-6}$. The induced constraint violations were practically negligible. $\Box$
Remark V.5 (OPF with constraint tightening).
When bounds on $\epsilon_{v}$, $\epsilon_{i}$, $E_{v}$, $E_{i}$, and the neural-network approximation error $\epsilon^{\textsf{NN}}$ are available or can be estimated numerically, the constraints of the original AC OPF problem (II-A) can be appropriately tightened so that the control policy generated by (8) renders the feasible set $\mathcal{S}$ forward invariant. In particular, leveraging the result of Proposition V.3, such a tightening guarantees that voltage and current limits are satisfied despite modeling, linearization, and approximation errors. This can be achieved by training the proposed NN-SGF using tightened voltage bounds given by $\underline{V}+\epsilon_{v}+2E_{v}+\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$ and $\overline{V}-\epsilon_{v}-2E_{v}-\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$, respectively.
We evaluated the performance of the NN-SGF under such tightened constraints. In the numerical experiments, we assumed negligible measurement noise (i.e., $\epsilon_{v}=0$), estimated the voltage linearization error bound as $E_{v}\approx 10^{-2}$ p.u., and obtained a neural-network approximation error of $\epsilon^{\textsf{NN}}=1.7\times 10^{-6}$. Substituting these values into Proposition V.3 yields tightened voltage limits of $0.9604$ and $1.0396$. The resulting AC OPF problem with tightened constraints remained feasible for all tested load and PV generation scenarios. The voltage profiles obtained in this experiment are reported in Fig. 6. As observed, the voltage magnitude constraints are never violated, demonstrating a robust mechanism to account for training and modeling errors. As expected, the resulting solution is conservative, owing to the bounding techniques employed in the derivation of Proposition V.3 (e.g., repeated use of triangle inequalities, as shown in the appendix). $\Box$
Figure 6: Numerical experiment with voltage constraints tightened based on the result of Proposition V.3. The NN-SGF is trained based on lower and upper voltage constraints given by $\underline{V}+\epsilon_{v}+2E_{v}+\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$ and $\overline{V}-\epsilon_{v}-2E_{v}-\beta^{-1}\|\boldsymbol{\Gamma}_{v}\|\epsilon^{\textsf{NN}}$, respectively. One can see that the voltage limits are satisfied strictly.
@@ VI Conclusions
We have proposed a solution method for solving the AC OPF where a neural network is used to approximate the solution of a convex QP defining the safe gradient flow. Our approach is shown to lead to both feedback-based online implementations and offline solutions based on power flow computations. Compared to existing methods that rely on neural networks, our algorithm ensures that the DERs’ setpoints are practically feasible and it ensures convergence to a neighborhood of a strict local optimizer of the AC OPF. These guarantees are important for power systems optimization tasks, as operating limits must be satisfied for a safe power delivery. Results are derived based on assumptions that are sufficient and conservative by design, as they rely on worst-case bounds to guarantee feasibility and forward invariance in the presence of modeling, measurement, and learning errors. In practice, the corresponding error bounds can be estimated empirically and were small in the numerical results. Future work will consider active set learning approaches for solving the QP associated with the SGF in the context of AC OPF in distribution grids [14], and will formalize results in terms of asymptotic stability under a discrete-time version of $\mathcal{F}^{\textsf{NN}}\!\big(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}\big)$, arising from, for example, forward-Euler methods. Finally, we will explore the derivation of formal results in terms of tracking and feasibility when the AC OPF problem is time-varying, under appropriate assumptions on the variability of the problem [38].
@@
Recall that $\boldsymbol{\nu}:=|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|+\boldsymbol{n}_{v}$, $\boldsymbol{\iota}:=|\boldsymbol{i}(\boldsymbol{u};\boldsymbol{s}_{l})|+\boldsymbol{n}_{i}$, $\boldsymbol{\xi}=(\boldsymbol{\nu},\boldsymbol{\iota})$; to streamline notation, we will use $|\boldsymbol{v}|$ and $|\boldsymbol{i}|$ to denote the error-free measurements or computations of voltage magnitudes and currents magnitudes. Recall that $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})=F_{\textsf{m}}(\boldsymbol{u},\boldsymbol{0},\boldsymbol{0},\boldsymbol{0})$. We express the NN-SGF controller as $\dot{\boldsymbol{u}}=\mathcal{F}^{\textsf{NN}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$, where $\boldsymbol{\theta}=(\boldsymbol{\theta}_{u,1},\ldots,\boldsymbol{\theta}_{u,G},\underline{V},\overline{V},\overline{I})$ contains the constraint parameters of the AC OPF. Rewrite the NN-SGF controller as:
@@
and is negative definite. This follows from properties of the safe gradient flow at strict local minimizers; see [1, Lemma 5.12] and [1, Proof of Theorem 5.7(iii)]. In particular, all eigenvalues of $\boldsymbol{J}_{F}$ are strictly negative, so $e_{1}:=-\lambda_{\max}(\boldsymbol{J}_{F})$ and $e_{2}:=-\lambda_{\min}(\boldsymbol{J}_{F})$ are well-defined positive real scalars.
Let $e_{1}:=-\lambda_{\max}(\boldsymbol{J}_{F})$, $e_{2}:=-\lambda_{\min}(\boldsymbol{J}_{F})$, and define the matrix
$P:=\int_{0}^{\infty}e^{\boldsymbol{J}_{F}^{\top}\zeta}e^{\boldsymbol{J}_{F}\zeta}\,d\zeta,$
@@
-B Proof of Proposition V.3
The proof leverages Nagumo’s Theorem. Consider the SGF controller $F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ in (11); for brevity, let $\hat{v}_{j}:=|\boldsymbol{\Gamma}_{v,j}\boldsymbol{u}+\bar{v}_{j}(\boldsymbol{s}_{l})|$ denote the linearized voltage magnitude at index $j\in\mathcal{M}$, and let the measurement noise $\boldsymbol{n}_{v}$ satisfy $\|\boldsymbol{n}_{v}\|_{\infty}\leq\epsilon_{v}$, as in Assumption 4. Moreover, by Assumption 3, the error between $|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|$ and its linear approximation is bounded by $E_{v}$. Thus, for every $j\in\mathcal{M}$,
$\hat{v}_{j}-E_{v}\leq|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\hat{v}_{j}+E_{v}.$ (18)
Next, from Assumption 4, we have $\|\boldsymbol{n}_{v}\|_{\infty}\leq\epsilon_{v}$. Applying triangle inequalities to the voltage with measurement noise gives,
$|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|-\epsilon_{v}\;\leq\;|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|+(\boldsymbol{n}_{v})_{j}\;\leq\;|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|+\epsilon_{v}.$ (19)
Recall that $\boldsymbol{\nu}_{j}=|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|+(\boldsymbol{n}_{v})_{j}$; therefore, we have the following:
$-\!\boldsymbol{\Gamma}_{v,j}^{\top}F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})\leq-\beta(\underline{V}-\boldsymbol{\nu}_{j})\leq-\beta\left(\underline{V}-|\hat{v}_{j}|-\epsilon_{v}-E_{v}\right),$
$\boldsymbol{\Gamma}_{v,j}^{\top}F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})\leq-\beta(\boldsymbol{\nu}_{j}-\overline{V})\leq-\beta\left(|\hat{v}_{j}|-(\overline{V}+\epsilon_{v}+E_{v})\right),$
where $E_{v}$ is defined in Assumption 3, and it is such that the error between each voltage magnitude and its linear approximation is bounded. Now consider the NN-SGF:
@@
$\leq-\beta(\underline{V}-|\hat{v}_{j}|-\epsilon_{v}-E_{v})+\|\boldsymbol{\Gamma}_{v,j}\|\epsilon^{\textsf{NN}}$
@@
$\leq-\beta(|\hat{v}_{j}|-\overline{V}-\epsilon_{v}-E_{v})+\|\boldsymbol{\Gamma}_{v,j}\|\epsilon^{\textsf{NN}}.$
To ensure the vector field is inward-pointing at the boundary of the voltage constraint set, we then require the linear approximation of the voltage magnitude at node $j\in\mathcal{M}$:
@@
$\mathcal{S}_{s,\hat{v}}$ $:=\left\{\boldsymbol{u}\in\mathcal{C}:\underline{V}_{s}\leq|\hat{v}_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V}_{s},j\in\mathcal{M}\right\},$
$\mathcal{S}_{s,\hat{\imath}}$ $:=\left\{\boldsymbol{u}\in\mathcal{C}:|\hat{i}_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I}_{s},j\in\mathcal{L}\right\},$
@@
Since the NN-SGF vector field is strictly inward-pointing, Nagumo’s Theorem implies that $\mathcal{S}_{s}$} is forward invariant under (8).
We note that the sets $\mathcal{S}_{s,\hat{v}}$ and $\mathcal{S}_{s,\hat{\imath}}$ are defined through inequalities involving the linearized voltage magnitude and linearized current magnitude, respectively. These inequalities, however, do not directly translate into constraints expressed in terms of the actual voltage and current magnitudes due to the nonlinear gap between the linearized and actual quantities.
To obtain a forward-invariant characterization in terms of the actual voltage and current magnitudes, we construct enlarged sets that contain $\mathcal{S}_{s}$ and are described by the actual quantities. To this end, observe that $|\hat{v}_{j}|-E_{v,j}\leq|v_{j}|\leq|\hat{v}_{j}|+E_{v,j},$ and $|\hat{\imath}_{j}|-E_{i,j}\leq|\imath_{j}|\leq|\hat{\imath}_{j}|+E_{i,j}.$ These bounds imply that $\mathcal{S}_{s,\hat{v}}\subseteq\mathcal{S}_{e,v}$ and $\mathcal{S}_{s,\hat{\imath}}\subseteq\mathcal{S}_{e,i},$ which further yields $\mathcal{S}_{s}\subseteq\mathcal{S}_{e},$ and the factor $2$ appearing in the definitions of $\underline{V}_{e}$, $\overline{V}_{e}$, and $\overline{I}_{e}$.